# and quotients of Coxeter groups

### v1

@@ 1. Introduction
In [BM], Barot and Marsh provided an explicit construction of presentation of a finite Weyl group $W$ for any quiver mutation-equivalent to an orientation of a Dynkin diagram with Weyl group $W$. In their construction, the Weyl group is presented as a quotient of a (usually infinite) Coxeter group by a normal closure of additional (cycle) relations. The result has already found some unexpected applications. It was shown in [FeTu2] that the normal closure of the cycle relations is torsion-free, which was used to construct geometric manifolds with the original Weyl group acting by isometries. A construction similar to the one in [BM] to obtain presentations of Artin braid groups for any quiver of finite type was provided in [GM].
The results of [BM] were extended in [FeTu1] to obtain presentations for all affine Weyl groups, where some new types of additional relations are required. Most quivers of finite and affine type can be realized as quivers of triangulations of marked bordered surfaces.
In the same paper, the construction was extended further to provide a group for every triangulation of an unpunctured marked bordered surface, where this group is invariant under change of triangulation and thus is presented as a quotient of numerous distinct Coxeter groups.
The aim of the present paper is to extend the construction to the case of quivers arising from punctured triangulated surfaces.
More precisely, we introduce a group for every triangulation of a marked surface not containing loops, and show that the group does not depend on the triangulation. The group is a quotient of a certain Coxeter group, and, when restricted to triangulations of unpunctured surfaces without loops, coincides with the one defined in [FeTu1] (see also [FeTu2, FeLSTu]).
@@ Acknowledgements
A.F. and P.T. are grateful to the Max Planck Institute for Mathematics in Bonn for the financial support and excellent research environment. M.S. thanks the Department of Mathematical Sciences of Durham University for support and excellent research atmosphere during his visit in the summer of 2024.
@@ 2.1. Quivers and mutations
±$\pm{c}\pm{d}={ab}$
Figure 2.1. Mutations of quivers. The sign before ${c}$ (resp., ${d}$) is positive if the three vertices form an oriented cycle, and negative otherwise. Either $c$ or $d$ may vanish. If $ab$ is equal to zero then neither value of $c$ nor orientation of the corresponding arrow does change.
@@ 2.2. Quivers of finite, affine, and finite mutation types
Let $\Sigma$ be a simply-laced Dynkin diagram (or a simply-laced affine Dynkin diagram different from $\widetilde{A}_{n}$). Then all orientations of $\Sigma$ are mutation-equivalent. The orientations of $\widetilde{A}_{n-1}$ split into $[n/2]$ mutation classes $\widetilde{A}_{k,n-k}$ (each class contains a cyclic representative with only two changes of orientations with $k$ consecutive arrows in one direction and $n-k$ in the other, $0<k<n$).
@@ 3.1. Construction
    $m_{ij}=\begin{cases}2&\text{if $i$ and $j$ are not joined;}\\ 3&\text{if $i$ and $j$ are joined by an arrow.}\end{cases}$
@@ 3.2. Invariance
()( ∗ ) $t_{i}=\begin{cases}s_{k}s_{i}s_{k}&\text{if there is an arrow $i\to k$ in $Q$,}\\ s_{i}&\text{otherwise},\end{cases}$
@@
Proposition 3.8 does not hold if we assume that $S$ has $3$ marked points only, an example of two distinct triangulations of thrice punctured genus $2$ surface is shown in Fig. 3.2. The triangulations are distinct indeed: cutting the surface along all arcs connecting points $b$ and $c$ in the triangulation on Fig. 3.2, left, results in the identification of sides of a $(2g+2)$-gon isomorphic to the original one, while cutting the surface along all arcs connecting points $b$ and $c$ in the triangulation on Fig. 3.2, right, results in an identification of sides different from the original one (more precisely, no pair of opposite sides is identified). None of the triangulations admits any loop-free flips.
Figure 3.2. Two triangulations of a thrice punctured surface of genus $2$. Sides colored in the same way are identified to get an oriented surface. Symmetries of the triangulation on the left act transitively on all three vertices, while the triangulation on the right has two different types of vertices.
@@ 4. Loop-free triangulations of closed surfaces
We proceed according to the following plan. In Section 4.1, we show that there always exists a loop-free triangulation of $S$, and describe a general form of such triangulation in the case of precisely three marked points. In Section 4.2, we show that using loop-free flips we can transform a triangulation to a specific form, namely, if $S$ is of genus $g$ with $m$ marked points, then the resulting triangulation is a triangulation of a genus $g$ surface with $3$ marked points and additionally with $m-3$ digons glued in along one of the arcs. In Sections 4.34.4, we prove Proposition 3.8 for surfaces with precisely four marked points, and in Section 4.5 we extend the proof to all closed surfaces with at least four marked points.
4.1. Surfaces with three marked points
@@
Denote by $g$ the genus of $S$. If $S$ has precisely three punctures, then $S$ can be glued from once punctured $(2g+2)$-gon $P$ by identifying opposite sides of $P$. Triangulating $P$ by arcs connecting the puncture with all vertices provides the required triangulation, see Fig. 4.1 (cf. Remark 3.9). To add a puncture on $S$, place it inside a triangle and connect with all three vertices.
Figure 4.1. Any loop-free triangulation of a closed surface of genus $g$ with three punctures can be obtained by a side-pairing of a punctured $(2g+2)$-gon triangulated as shown in the figure.
@@
It is easy to see that all loop-free triangulations of surfaces with $3$ marked points can be constructed as in Lemma 4.1 with different identifications of sides of the $(2g+2)$-gon.
@@
4.2. Reducing the number of marked points to $3$
@@
Let $S$ be a closed surface with a loop-free triangulation $T$ as in Remark 4.11, and let $\tilde{T}$ be the associate triangulation of $\tilde{S}$. Let $\tilde{f}$ be a loop-free flip of $\tilde{T}$. Then there exists a sequence of loop-free flips taking $T$ to a triangulation $T^{\prime}$ such that $\tilde{T}^{\prime}$ coincides with $\tilde{f}(\tilde{T})$.
@@
Let $\tilde{\gamma}$ be the arc of $\tilde{T}$ changed by flip $\tilde{f}$. Then $\pi^{-1}(\tilde{\gamma})$ is either an arc of $T$ or a union of digons. If $\pi^{-1}(\tilde{\gamma})$ is an arc, then the flip in this arc produces the required triangulation $T^{\prime}$. Otherwise, due to Remark 4.12, we can apply Lemma 4.10(c) to create a triangulation with the same associate triangulation $\tilde{T}$ but with $\pi^{-1}(\tilde{\gamma})$ being a single arc.
@@ 4.3. Regluing polygons
Let $P$ be a $(2g+2)$-gon with an identification of sides resulting in a twice punctured surface $S$ of genus $g$, such that every side of $P$ has two distinct endpoints. Define an admissible regluing of $P$ as the following two-step procedure.
@@
The result of the procedure is again a $(2g+2)$-gon with another identification of sides resulting in the same surface $S$.
@@
Let $P$ be a $(2g+2)$-gon with an identification of sides resulting in a twice punctured surface of genus $g$, such that every side of $P$ has two distinct endpoints. There exists a sequence of admissible regluings resulting in a polygon in which all pairs of opposite sides are identified.
@@
Label the sides of $P$ by $e_{1},\dots,e_{g+1}$, where the identified sides have the same label. If we read the labels clockwise in the target polygon, we get the same word repeated twice. Now read the labels clockwise in $P$, they produce a word $w$. We prove the following claim: if $W$ contains a word $e_{i_{1}}\dots e_{i_{k}}$ and also a separate word $e_{i_{2}}\dots e_{i_{k}}$, then there exists an admissible regluing such that the resulting word contains two copies of a word of length $k$. Applying the claim repeatedly, we obtain the statement of the lemma.
@@
Figure 4.4. To the proof of Lemma 4.15: the case of the second side labeled $e_{i_{1}}$ lying between $e_{i_{k}}$ and $e_{i_{2}}$ in clockwise order.
@@ 4.4. Surfaces with four marked points
Let $P$ be a once punctured $(2g+2)$-gon with an identification of sides resulting in a thrice punctured surface $\tilde{S}$ of genus $g$, such that every side of $P$ has two distinct endpoints. Let $\tilde{T}$ be a triangulation of $\tilde{S}$ containing all sides of $P$ and all arcs connecting the puncture inside $P$ with vertices of $P$. Consider a surface $S$ with four marked points and a loop-free triangulation $T$ containing a digon, such that the associate triangulation of $T$ is precisely $\tilde{T}$.
@@
Let $(P_{1},T_{1})$ and $(P_{2},T_{2})$ be two pairs satisfying the assumptions of Lemma 4.16, i.e. both $P_{1}$ and $P_{2}$ are $(2g+2)$-gons, but the identifications of sides are different. Assume that $P_{2}$ can be obtained from $P_{1}$ by an admissible regluing. Then $T_{2}$ can be obtained from $T_{1}$ by a sequence of loop-free flips.
@@ 5.2. Bordered surfaces with three marked points
Figure 5.5. Constructing a loop-free triangulation of a bordered surface with three marked points from a triangulation of a closed surface.
As for closed surfaces, all loop-free triangulations of surfaces with $3$ marked points can be constructed as in Lemma 5.5 with different identification of sides of the $(2g+2)$-gon.
@@ 6.1. Triangulations with self-folded triangles
[…]
@@ 6.2. Orbifolds
    $m_{ij}=\begin{cases}2&\text{if $i$ and $j$ are not joined;}\\ 3&\text{if $i$ and $j$ are joined by a simple arrow;}\\ 4&\text{if $i$ and $j$ are joined by an arrow of weight $2$};\end{cases}$
@@
[…]
@@ 6.5. Example
and $N(Q)$ is the normal closure of $(t_{1}t_{4}t_{3}t_{5}t_{3}t_{4})^{2}$ in $\tilde{G}(Q)$, see Remark 3.2.
@@
In fact, this homomorphism is not an isomorphism: the kernel is the normal closure of the element $(t_{3}t_{2}t_{1}t_{4}t_{1}t_{2})^{2}\in G(Q)$ (we use MAGMA [BCP] to verify this).

### v2

@@ 1. Introduction
In [BM], Barot and Marsh provided an explicit construction of presentation of a finite Weyl group $W$ for any quiver mutation-equivalent to an orientation of a Dynkin diagram with Weyl group $W$. In their construction, the Weyl group is presented as a quotient of a (usually infinite) Coxeter group by a normal closure of additional (cycle) relations. The result has already found some unexpected applications. It was shown in [FeTu2] that the normal closure of the cycle relations is torsion-free, which was used to construct geometric manifolds with the original Weyl group acting by isometries.
The results of [BM] were extended in [FeTu1] to obtain presentations for all affine Weyl groups, where some new types of additional relations are required. Most quivers of finite and affine type can be realized as quivers of triangulations of marked bordered surfaces. In the same paper, the construction was extended further to provide a group for every triangulation of an unpunctured marked bordered surface, where this group is invariant under change of triangulation and thus is presented as a quotient of numerous distinct Coxeter groups.
A construction similar to the one in [BM] to obtain presentations of Artin braid groups for any quiver of finite type (with potential) was provided by Grant and Marsh in [GM] (see also [Qy1]). These results were generalized by Qiu and Zhou in [QZ], where a finite presentation for a braid group for any unpunctured surface was constructed (and it was shown that this group coincides with the braid twist group). Taking the quotient of this braid group by setting all generators to be involutions recovers the group constructed in [FeTu1].
The results of [QZ] were generalized to punctured surfaces with boundary by Qiu [Qy2]. More precisely, a finite presentation of an invariantly defined mixed twist group was given in any triangulation with all punctures contained in self-folded triangles.
The aim of the present paper is to extend the construction from [FeTu1] to the case of quivers arising from punctured triangulated surfaces.
More precisely, we introduce a group for every triangulation of a marked surface not containing loops, and show that the group does not depend on the triangulation. The group is a quotient of a certain Coxeter group, and, when restricted to triangulations of unpunctured surfaces without loops, coincides with the one defined in [FeTu1] (see also [FeTu2, FeLSTu]). The construction allows us to extend the results of [FeLSTu] to punctured surfaces by establishing connections with extended affine Weyl groups, this will appear in a forthcoming paper.
@@
For surfaces with boundary, the group $G=G(Q)$ turns out to be a finite index subgroup of the quotient of the mixed twist group from [Qy2]. To obtain $G$, one needs to consider a triangulation of a particular type (see Section 6.5) and impose the relations on the mixed twist group by setting all the generators (in the presentation given in [Qy2]) to be involutions, and then take a subgroup of index $2^{p}$, where $p$ is the number of punctures. We note that imposing additional relations allows us to give presentation of our group in more general settings, see Section 6.5 for the details.
@@ Acknowledgements
We thank the anonymous referee for pointing out the connections to results of [Qy1, QZ, Qy2]. A.F. and P.T. are grateful to the Max Planck Institute for Mathematics in Bonn for the financial support and excellent research environment. M.S. thanks the Department of Mathematical Sciences of Durham University for support and excellent research atmosphere during his visit in the summer of 2024.
@@ 2.1. Quivers and mutations
$\pm{c}\pm{d}={ab}$
Figure 2.1. Mutations of quivers. The sign before c{c} (resp., d{d}) is positive if the three vertices form an oriented cycle, and negative otherwise. Either $c$ or $d$ may vanish. If $ab$ is equal to zero then neither value of $c$ nor orientation of the corresponding arrow does change.
@@ 2.2. Quivers of finite, affine, and finite mutation types
Let $\Sigma$ be a simply-laced Dynkin diagram (or a simply-laced affine Dynkin diagram different from $\widetilde{A}_{n}$). Then all orientations of $\Sigma$ are mutation-equivalent. The orientations of $\widetilde{A}_{n-1}$ split into $[n/2]$ mutation classes $\widetilde{A}_{k,n-k}$ (each class contains a cyclic representative with only two changes of orientations with $k$ consecutive arrows in one direction and nkn-k in the other, $0<k<n$).
@@ 3.1. Construction
[…]
@@ 3.2. Invariance
() $t_{i}=\begin{cases}s_{k}s_{i}s_{k}&\text{if there is an arrow $i\to k$ in $Q$,}\\ s_{i}&\text{otherwise},\end{cases}$
@@
Proposition 3.8 does not hold if we assume that $S$ has $3$ marked points only, an example of two distinct triangulations of thrice punctured genus $2$ surface is shown in Fig. 3.2. The triangulations are distinct indeed: cutting the surface along all arcs connecting points $b$ and $c$ in the triangulation on Fig. 3.2, left, results in the identification of sides of a $(4g+2)$-gon isomorphic to the original one, while cutting the surface along all arcs connecting points $b$ and $c$ in the triangulation on Fig. 3.2, right, results in an identification of sides different from the original one (more precisely, no pair of opposite sides is identified). None of the triangulations admits any loop-free flips.
Figure 3.2. Two triangulations of a thrice punctured surface of genus 22. Sides colored in the same way are identified to get an oriented surface. Symmetries of the triangulation on the left act transitively on all three vertices, while the triangulation on the right has two different types of vertices.
@@ 4. Loop-free triangulations of closed surfaces
We proceed according to the following plan. In Section 4.1, we show that there always exists a loop-free triangulation of $S$, and describe a general form of such triangulation in the case of precisely three marked points. In Section 4.2, we show that using loop-free flips we can transform a triangulation to a specific form, namely, if $S$ is of genus $g$ with $m$ marked points, then the resulting triangulation is a triangulation of a genus $g$ surface with $3$ marked points and additionally with m3m-3 digons glued in along one of the arcs. In Sections 4.34.4, we prove Proposition 3.8 for surfaces with precisely four marked points, and in Section 4.5 we extend the proof to all closed surfaces with at least four marked points.
4.1. Closed surfaces with three marked points
@@
Denote by $g$ the genus of $S$. If $S$ has precisely three punctures, then $S$ can be glued from once punctured $(4g+2)$-gon $P$ by identifying opposite sides of $P$. Triangulating $P$ by arcs connecting the puncture with all vertices provides the required triangulation, see Fig. 4.1 (cf. Remark 3.9). To add a puncture on $S$, place it inside a triangle and connect with all three vertices.
Figure 4.1. Any loop-free triangulation of a closed surface of genus $g$ with three punctures can be obtained by a side-pairing of a punctured $(4g+2)$-gon triangulated as shown in the figure.
@@
It is easy to see that all loop-free triangulations of surfaces with $3$ marked points can be constructed as in Lemma 4.1 with different identifications of sides of the $(4g+2)$-gon.
@@
4.2. Reducing the number of marked points to 33
@@
[…]
@@
[…]
@@ 4.3. Regluing polygons
Let $P$ be a $(4g+2)$-gon with an identification of sides resulting in a twice punctured surface $S$ of genus $g$, such that every side of $P$ has two distinct endpoints. Define an admissible regluing of $P$ as the following two-step procedure.
@@
The result of the procedure is again a $(4g+2)$-gon with another identification of sides resulting in the same surface $S$.
@@
Let $P$ be a $(4g+2)$-gon with an identification of sides resulting in a twice punctured surface of genus $g$, such that every side of $P$ has two distinct endpoints. There exists a sequence of admissible regluings resulting in a polygon in which all pairs of opposite sides are identified.
@@
Label the sides of $P$ by $e_{1},\dots,e_{2g+1}$, where the identified sides have the same label. If we read the labels clockwise in the target polygon, we get the same word repeated twice. Now read the labels clockwise in $P$, they produce a word $w$. We prove the following claim: if $W$ contains a word $e_{i_{1}}\dots e_{i_{k}}$ and also a separate word $e_{i_{2}}\dots e_{i_{k}}$, then there exists an admissible regluing such that the resulting word contains two copies of a word of length $k$. Applying the claim repeatedly, we obtain the statement of the lemma.
@@
Figure $4.$ To the proof of Lemma 4.15: the case of the second side labeled $e_{i_{1}}$ lying between $e_{i_{k}}$ and $e_{i_{2}}$ in clockwise order.
@@ 4.4. Surfaces with four marked points
Let $P$ be a once punctured $(4g+2)$-gon with an identification of sides resulting in a thrice punctured surface $\tilde{S}$ of genus $g$, such that every side of $P$ has two distinct endpoints. Let $\tilde{T}$ be a triangulation of $\tilde{S}$ containing all sides of $P$ and all arcs connecting the puncture inside $P$ with vertices of $P$. Consider a surface $S$ with four marked points and a loop-free triangulation $T$ containing a digon, such that the associate triangulation of $T$ is precisely $\tilde{T}$.
@@
Let $(P_{1},T_{1})$ and $(P_{2},T_{2})$ be two pairs satisfying the assumptions of Lemma 4.16, i.e. both $P_{1}$ and $P_{2}$ are $(4g+2)$-gons, but the identifications of sides are different. Assume that $P_{2}$ can be obtained from $P_{1}$ by an admissible regluing. Then $T_{2}$ can be obtained from $T_{1}$ by a sequence of loop-free flips.
@@ 5.2. Bordered surfaces with three marked points
Figure $5.$ Constructing a loop-free triangulation of a bordered surface with three marked points from a triangulation of a closed surface.
As for closed surfaces, all loop-free triangulations of surfaces with $3$ marked points can be constructed as in Lemma 5.5 with different identification of sides of the $(4g+2)$-gon.
@@ 6.1. Triangulations with self-folded triangles
Let $S$ be a marked surface with at least $4$ features, let $T$ be a triangulation of $S$ not containing loops, and let $Q$ be the corresponding quiver. Then the group $G(Q)$ with defining relations $(\mathrm{R}1)$$(\mathrm{R}4)$ is an invariant of $S$, i.e. it does not depend on the triangulation.
@@ 6.2. Orbifolds
[…]
@@
Let $S$ be a marked orbifold with orbifold points of order $2$, suppose that $S$ has at least $4$ features. Let $T$ be a triangulation of $S$ not containing loops, and let $\Sigma$} be the corresponding diagram. Then the group $G(\Sigma)$ with defining relations $(\mathrm{R}1^{\prime})$$(\mathrm{R}4^{\prime})$ is an invariant of $S$, i.e. it does not depend on the triangulation.
@@ 6.4.2. Three features
6.5. Relation to mixed twist groups
For a marked surface with non-empty boundary, Qiu [Qy2, Section 2.5] defined a mixed twist group $\mathrm{MT}$ of the surface and provided its finite presentation for any (ideal) triangulation $T$ in which every puncture is contained in a self-folded triangle. If the corresponding tagged triangulation contains no loops, the presentation of $\mathrm{MT}$ is simplified to the one shown in [Qy2, Corollary 2.13], and our group $G(Q)$ can be obtained from $\mathrm{MT}$ as follows.
The group $\mathrm{MT}$ is generated by $L$-twists (corresponding to self-folded arcs of $T$) and braid twists (corresponding to other arcs of $T$), with relations being either triangle relations or higher braid relations. First, we set all generators of $\mathrm{MT}$ to be involutions. Next, for every self-folded triangle we take the corresponding $L$-twist $a$ and braid twist $b$, and substitute the generator $a$ with abaaba (this now corresponds to a tagged arc in the tagged triangulation), which is equivalent to taking a subgroup of index $2^{p}$, where $p$ is the number of punctures. It is easy to see that higher braid relations in the presentation of $\mathrm{MT}$ become Coxeter relations (R2). The triangle relations now become cycle relations (R3) by [GM, Lemma 2.5], so we obtain precisely the group $G(Q)$.
Since the group $\mathrm{MT}$ is defined invariantly, we can use this correspondence to extend our results.
6.5.1. Triangulations with all punctures contained in self-folded triangles
If all punctures in a (tagged) triangulation are contained in self-folded triangles, then the corresponding quiver $Q$ does not contain oriented cycles of length greater than $3$. For such triangulations (even with loops), we can also write down the presentation of $G(Q)$ following the rules from [FeTu1]. This can be checked explicitly, or deduced from [Qy2]. Note that such triangulations exist for surfaces with non-empty boundary only.
6.5.2. Surfaces with boundary
The relation to the mixed twist group provides another proof of the invariance of the group $G(Q)$ for surfaces with non-empty boundary. Moreover, due to the section above, we can define the group for any marked surface with boundary, thus dropping the requirement to have at least four features.
@@ 6.6. Example
and $N(Q)$ is the normal closure of $(t_{1}t_{2}t_{3}t_{4}t_{3}t_{2})^{2}$ in $\tilde{G}(Q)$, see Remark 3.2.
@@
In fact, this homomorphism is not an isomorphism: the kernel is the normal closure of the element $(t_{1}t_{5}t_{3}t_{4}t_{3}t_{5})^{2}\in G(Q)$ (we use MAGMA [BCP] to verify this).